FLOSS01: Logbook Entries

FLOSS01: Site all stns Messages, 35 Entries..

Return to Logbook Contents Page
Entry Date Title Site Author #Graphics
178 Wed 27-Mar-2002Radiometers and soil sensors removed from all sitesall stnsdelany
173 Wed 13-Mar-2002rad & kryptons cleanedall stnsoncley
153 Fri 22-Feb-2002Kr and Rad cleanall stnsdelany
148 Thu 21-Feb-2002Radiometer visitsall stnsdelany
146 Wed 20-Feb-2002Soil samples at sage, bare and towerall stnsdelany
139 Mon 18-Feb-2002kryptons cleanedall stnsoncley
138 Mon 18-Feb-2002rads cleanedall stnsoncley
137 Mon 18-Feb-2002power system inventoryall stnsoncley
133 Tue 12-Feb-2002all stnsdelany
118 Thu 17-Jan-2002soil samplesall stnssemmer
116 Thu 17-Jan-2002soil sample from Tony's visitall stnssemmer
111 Mon 07-Jan-2002Visit 01/04/2002 Task CheckList Reviewall stnsmilitzer
106 Fri 04-Jan-2002Radiometer Domes cleanedall stnsmilitzer
100 Wed 26-Dec-2001temperature event?all stnsoncley
98 Wed 26-Dec-2001temp sensitivitiesall stnsoncley
96 Wed 19-Dec-2001archive size and disk capacityall stnsmaclean
89 Thu 13-Dec-2001calibration files for CR10Xsall stnssemmer
82 Sun 09-Dec-2001rebooted asterall stnsoncley
79 Sun 09-Dec-2001visitorsall stnsoncley
76 Sat 08-Dec-2001shallow inversion todayall stnsoncley
72 Sat 08-Dec-2001TRH 703all stnsoncley
70 Sat 08-Dec-2001serial numbers and heightsall stnsoncley
59 Thu 06-Dec-2001radiometers and soil sensorsall stnsdelany1
44 Tue 04-Dec-2001Pressure errorall stnsoncley1
43 Tue 04-Dec-2001soil moistureall stnsoncley
32 Mon 03-Dec-2001task listall stnsoncley
29 Mon 03-Dec-2001daily statusall stnsoncley
26 Sun 02-Dec-2001data network configurationall stnsmaclean
25 Sun 02-Dec-2001Dec 2 morning statusall stnsmaclean
22 Sat 01-Dec-2001Operational!all stnsmaclean
15 Sat 05-Aug-2000Soil water gravimetric dataall stnsdelany
14 Thu 29-Nov-2001statusall stnsoncley
8 Tue 20-Nov-2001manual control noteall stnssemmer
5 Sat 10-Nov-2001Breezecom settingsall stnsmaclean
4 Sat 10-Nov-2001Adam/network testingall stnsmaclean


178: Radiometers, Site all stns, Wed 27-Mar-2002 09:41:17 MST, Radiometers and soil sensors removed from all sites
A special trip was made to FLOSS to recover the radiometers and soil probes
Tony, Charlie and Pierce removed the sensors at Bare, Sage and then Tower

(Also the EVE boxes were removed at the bare and sage sites - so no
data from those sites after 15:00 MST March 26.   GDM)

Bare 12:15
Ground not frozen and soil sensors easily uprooted
Net Q97035
Fluxplate H993563
Tsoil NCAR 0001
Soil moisture 013
Soil sensors placed in Netrad box

Campbell 13527

Sage 12:30 
The up looking Pyrg and K&Z were each shadowed using Steve O's wand, for
three successive periods of 30 s at ~ 12:40
Net Q94125
Flux
SoilT
SoilQ

Campbell 11173

Tower 13:30
The up looking Pyrg and K&Z were each shadowed using Steve O's wand, for
three succesive 30 s periods, for three 10 s periods and for one 60s period


Net Q99258
Flux
SoilT
SoilQ

Campbell 11174

The CO2 cal gases: 
#6 Local 0/0
#5 Hical 700/20


173: Kryptons, Site all stns, Wed 13-Mar-2002 16:51:47 MST, rad & kryptons cleaned
Sites 2 first then 3.

153: Kryptons, Site all stns, Fri 22-Feb-2002 11:25:17 MST, Kr and Rad clean
At 10:50 visited Sage and cleaned Kr and all radiometers
At 11:00 visited Bare and cleaned Kr and netrad
At 11:10 visited tower and cleaned Kr.2m and all radiometers

Note in no case was there any indication of dirt on the domes.
148: Radiometers, Site all stns, Thu 21-Feb-2002 12:38:40 MST, Radiometer visits
At 09:45 visited Sage. No frost on radiometers but some on frost detector and 
on top dome of net radiometer. Cleaned frost from frost detector and net 
radiometer. Handkerchief placed over frost detector for 10 minutes.
All fans operating. Power dissapation resister cold.

At 10:08 visited tower site. No frost on radiometers nor frost detector. Net
radiometer had frost on upper dome. Covered uplooking frost detector from 10:09
until 10:30. Covered downlooking frost detector from 10:30 til 10:40.

146: Soil, Site all stns, Wed 20-Feb-2002 16:30:14 MST, Soil samples at sage, bare and tower
At 10:00 0-5cm soil core taken at sage. Several cm of fresh snow first removed.
      Fresh    Dry
AI   19.4g    16.5g
AII  20.4g    17.4g
AIII 23.4g    19.7g

At 10:30 0-5cm soil core taken at bare. Several cm of fresh snow first removed.
      Fresh    Dry
BI   19.6g    16.7g
BII  23.9g    20.1g
BIII 27.1g    22.7g

At 11:00 0-5cm soil core taken at tower A cm of fresh snow first removed.
      Fresh    Dry
CI   17.6g    13.2g
CII  17.0g    12.7g
CIII 16.6g    12.0g

The soil samples were taken  on Wed, 20 Feb''02, and weighed
They were cooked in the oven at 110C and reweighed on Thu. To check the 
completeness of drying they were left in the oven until Fri AM and reweighed.
The Thu and Fri weights were identical.
139: Kryptons, Site all stns, Mon 18-Feb-2002 15:47:14 MST, kryptons cleaned
All kryptons have now been cleaned, starting with the top of the main tower at
about 13:00 and progressing down, the sage ~1500 and bare ~1530.

138: Radiometers, Site all stns, Mon 18-Feb-2002 15:46:12 MST, rads cleaned
All rads have now been cleaned.
13:55 local - main tower
~15:00 local sage site
~15:15 local bare site

137: Power, Site all stns, Mon 18-Feb-2002 15:43:45 MST, power system inventory
PAM charge controllers (inventory for IHOP):

bare:
main box: Flexcharge NC25A-12 and Sunsaver 20 in cooler + 5 batteries
fence charger: Sunsaver 20 in old-style box + 1 battery


sage:
main box: Flexcharge NC25A-12 and Sunsaver 20 in cooler + 4 batteries
fence charger: Prostar 30 in old-style box + 1 battery

133: Status, Site all stns, Tue 12-Feb-2002 14:44:22 MST,
FLOSS status, 2002 Feb 11

NOTE: I have just started the data minding task


Webcam:


        Meteorology
        -----------
        Prop Spd/ Dir:  ok     
        T/RH (ISFF):    ok
        Pressure:       ok
        Rainr:



        Depth:          ok value varies
        Rsw:            ok the albedo at the tower is ~60%, at sage only 30%
        Rlw:            ok
        Rnet:           ok different albedos

        Soil
        ----

        Tsoil:          ok, tower still noticably warmer
        Gsoil:          ok, sage still has significantly less flux
        Qsoil:          ok, sage still driest


                        Sonics
        ------
        flags/diag              ok, flag and diag hits on tower 2m.  This appears
                        correlated with temps below -20degC.
        spd, dir, w:            ok the difference plot now fixed        
        variances:              ok the differnce plot now fixed
        u*, sigma_w/u*, w'tc':  ok


        Scalars
        -------

        H2O             ok, The open path Licor is strange beast, probably a calib.
                        issue.  It apparently has inverse trend and
                        gain factor of about 12x compared with kryptons.
                        Sage/Bare still have irregularly high signals
                        at times and these correlate with very low temps
                        going below ~-22degC. 
        co2:            ok still showing calibration peaks, ie still got cal gas
        w'h2o':         ok
        w'co2':         ok


        Status
        ------
        PAM power
        X/Ylevel
        Ice    

        <><><><><><><><><><><><><><><><><><><><><><><><><><><><><><><><><>



118: Soil, Site all stns, Thu 17-Jan-2002 17:44:18 MST, soil samples
soil samples:  

wet
	tower		sage
	a1-14.3		b1-29.7
	a2-16.3		b2-31.0
	a3-16.3		b3-31.3
dry (weighed 2/18/02 by spo)
        a1-15.9		b1-25.7
	a2-17.4		b2-26.4
	a3-18.0		b3-27.0
(how can the an samples be higher???)


116: Soil, Site all stns, Thu 17-Jan-2002 10:56:41 MST, soil sample from Tony's visit
These are the dry numbers. Could not find enter for
the wet numbers

A1 - 5.8
A2 - 7.1  
A3 - 9.0

B1 - 16.0
B2 - 15.0
B3 - 15.8
  
111: Logistics, Site all stns, Mon 07-Jan-2002 10:33:12 MST, Visit 01/04/2002 Task CheckList Review
====================================================================
			FLOSS Service Visit Tasks
				4-Jan.-02
====================================================================

TASKS:
Done: - Site First:
	Place stool under Snow Depth Gauge SR50 while there.
	Later on checked and exposed thermocouple to atmosphere.

Done: - Tower Guy Wire Tension check
	(update logbook entry 55)
	Confirm they are consistent before climbing...
	(alert Kurt if change by more than 50 lbs)

Done: - Tower Thermocouples
	Install 6-each Thermocouples
	Retrieve dead Thermocouples on lower levels

Done: - Tower Photographs
	near top of tower
	See sample photographs with camera.

Done: - Tower 20m Prop Change/Fix dead

Done: - Tower Clean Kryptons	 (alcohol / Q-tips)

Done: - Tower	Clean Radiometers (dry wipes)

Done: - Tower Cal-Gas
	Change 2 cylinders  (update logbook entry 33)
	check flow rate at CO2 sensor inlet
NOT: 	(create logbook entry on flow rate)
NOT:	check	turn on/off:	'ttl_on cosmos 1'
				'ttl_on cosmos 0'

Done: - Tower Precip Gauge fluid level check
	(drain if within 6" top of the gauge)

Done: - PAMs	Clean Kryptons	 (alcohol / Q-tips)
Done: - PAMs	Clean Radiometers (dry wipes)
NOT:  - PAMs	Find/Retrieve Prototype EVE Power Supply
	Could not find this power supply...bummer...
NOT:  - PAM	Bare site Vbatt monitor bad
	Take replacement bbox next time

NOT:  - Trime soil moisture reading; each site
	(update logbook entry 63)
        Battery dead; recharge and take next time.

Done: - Gravimetric soil sample at each site
	(logbook entry 52; clone logbook entry 62)

Done: - Base	Swap backup disk in aster
	Reboot / Check aster

Done: - Base	Retrieve Licor Manual

Done: - Base	Update LOGBOOK

Done: *** DON'T FORGET: REMOVE STOOL FROM UNDER SR50 ***





106: Radiometers, Site all stns, Fri 04-Jan-2002 15:57:35 MST, Radiometer Domes cleaned
12:45 Radiometers on Tower, all 6 fans OK, Rnet level OK.

15:30 Radiometers on Sage-brush site, all fans OK, Rnet level OK,

On both sites the domes appeared clean before wiping off.

On both sites there was some snow on top of the up-looking radiometers shield.
100: Temp_RH, Site all stns, Wed 26-Dec-2001 12:13:04 MST, temperature event?
For a few hours before dawn this morning, the bare and sage temperatures were
several degrees colder than at the main tower.  The only quick explanation I
can come up with is that the main tower was in a different air mass than the
other two stations.

98: , Site all stns, Wed 26-Dec-2001 11:55:35 MST, temp sensitivities
I now suspect temperature sentivities of several sensors:
- CO2 closed-path
- CO2 open-path
- (ASTER?) Barometer
- Snow depth gauge

All should be characterized in some fashion:
- CO2 closed-path by processing cal gas values
- CO2 open-path by comparison to corrected closed-path???
- Barometer through cal chamber check (post cal)
- Snow depth through "stool" check (also post cal in chamber)

96: Base, Site all stns, Wed 19-Dec-2001 10:54:00 MST, archive size and disk capacity
Archive file sizes (Kbyte = 1024 byte) per day:

	marigold: 209396
	cosmos:   177885
	s2:        36642
	s3:        36987
        aster:       185

	total:    461095

The /data disk is: 11442156 Kbytes, so we can store 24 days on /data.

Backup disk #2 is 16547932 Kbytes, which holds 35 days.

Backup disk #1 is slightly larger than #2, something like 18 Gbytes.

Note that these sizes don't include any photos.

89: Radiometers, Site all stns, Thu 13-Dec-2001 10:45:06 MST, calibration files for CR10Xs
Below yare the calibration coefficients used in
each CR10X.
NOTE: Before the installation of stn3 at the "sage" site, stn2 at "bare"
had the 4 comp system. This system was moved to stn3.
The soil sensors and the Q7 stayed at stn2 and a new
set of soil sensors and Q7 along with the 4 comp were
installed at stn3. So if you are looking at data for
stn2 before the installation of stn3, you need to use
the 4 comp coefficients listed in stn3.

TOWER

sensor	SN	     location	calibration
Q7	Q99258	A+	00	9.35
		A-	01	11.55
PYGIN	26416		02	280.1120448
PYGOUT	29260		03	311.5264798
KZIN	940187		04	100.7049345
KZOUT	940186		05	82.98755187
HFT	H983212		06	38.4
PIcase			10	2500
		A0	11	0.00335513
		A1	12	0.278704
		A2	13	1.4317
		A3	14	998.718
PIdome			15	2500
		A0	16	0.00305339
		A1	17	0.268816
		A2	18	8.77845
		A3	19	-2267.84
POcase			20	2500
		A0	21	0.00335427
		A1	22	0.279319
		A2	23	5.19638
		A3	24	-247.625
POdome			25	2500
		A1	26	0.00305418
		A2	27	0.268098
		A3	28	4.06691
		A4	29	403.649
Tsoil	NCAR002	A0	30	-241.477805
		A1	31	231.052081
		A2	32	9.262642
level  N0383 X, A0	33	0.05750311
		A1	34	14.26792000
		A2	35	-0.13395810
		A3	36	-0.20585990
		A4	37	0.0000000
		A5	38	0.0000000
	     Y, A0	39	-0.02092993
		A1	40	14.44724000
		A2	41	0.12732100
		A3	42	-0.39985070
		A4	43	0.0000000
		A5	44	0.0000000
CS615	12	A0	45	-0.187
		A1	46	0.037
		A2	47	0.335
ice on/off 1=on, 0=off	50	1
min bat	(volts)		51	11
X time	(secs)		52	300
outmode	0=a, 1=b	55	0


stn2 at bare site
sensor	SN	     location	calibration	
Q7	Q97035	A+	00	9.26	
		A-	01	11.52	
HFT	H993563		06	38.9	
Tsoil	NCAR001	A0	30	-241.6249	
		A1	31	230.9534	
		A2	32	9.0362	
CS615	13	A0	45	-0.187	
		A1	46	0.037	
		A2	47	0.335	

stn3 at sage site
sensor	SN	     location	calibration   (sensors at stn2 before stn3)
Q7	Q94125	A+	00	8.64			
		A-	01	10.78
PYGIN	31978		02	280.8594299		*
PYGOUT	31976		03	259.6053998		*
KZIN	940185		04	80.25682183		*
KZOUT	970380		05	86.50519031		*
HFT	H993560		06	37.7
PIcase			10	2500			*
		A0	11	0.00335405		*
		A1	12	0.28145			*
		A2	13	3.04961			*
		A3	14	289.824			*
PIdome			15	2500			*
		A0	16	0.0030332		*
		A1	17	0.291681		*
		A2	18	-3.4779			*
		A3	19	870.025			*
POcase			20	2500			*
		A0	21	0.00335116		*
		A1	22	0.285883		*
		A2	23	0.220836		*
		A3	24	756.272			*
POdome			25	2500			*
		A1	26	0.00302872		*
		A2	27	0.297063		*
		A3	28	-5.33875		*
		A4	29	1070.53			*
Tsoil	NCAR003	A0	30	-242.721555
		A1	31	233.953195
		A2	32	8.282468
level  N2018 X, A0	33	-0.3555644		*
		A1	34	11.0103000		*
		A2	35	-0.1015489		*
		A3	36	-0.4473981		*
		A4	37	0.0000000		*
		A5	38	0.0000000		*
	     Y, A0	39	0.5774926		*
		A1	40	10.6420400		*
		A2	41	0.0040844		*
		A3	42	-0.1290564		*
		A4	43	0.0000000		*
		A5	44	0.0000000		*
CS615	8	A0	45	-0.187
		A1	46	0.037
		A2	47	0.335
ice on/off 1=on, 0=off	50	0
min bat	(volts)		51	0
X time	(secs)		52	0
outmode	0=a, 0=b	55	1



82: Base, Site all stns, Sun 09-Dec-2001 18:26:41 MST, rebooted aster
The reason why Xlevel and Ylevel have been bad is that channel config had too
few values for the logger at the bare site (cr10_13, rather than cr10_18).
Since I don't know how to restart ingest for the PAM sites (I tried
ingest_control -r stn2, but this didn't work) and I couldn't reach Gordon,
I simply restarted aster.  This appears to have worked (though I still had
to start cockpit stn2 manually).

Now Xlevel and Ylevel.bare appear to be okay.

The biggest problem with the way we were collecting data is that we dropped
all the Qsoilraw.bare values.  Thus, it will be extra work to revise the 
calibrations for the Qsoil.bare data taken to date.

79: Status, Site all stns, Sun 09-Dec-2001 13:28:38 MST, visitors
Bob Kelly (Univ. of Wyoming) was here for a tour this morning.

Other visitors we've had:

Ed - (the local NWS cooperative observer for the past 25 years) was here for
a tour Friday afternoon.

Todd S- and Pam - from the Wildlife Refuge were here Thursday to teach Larry
and Dean about local plants.

Mark Lanier from the Wildlife Refuge was here Wed? last week to see what we
were up to.

We've (mostly Larry) also fielded reporter phone calls from the Denver Post 
and Colorado Springs Gazette last week.

We're expecting a visit by Matt Landis and other local High School teachers
tomorrow afternoon.

76: Status, Site all stns, Sat 08-Dec-2001 17:29:35 MST, shallow inversion today
We noticed a very shallow inversion this afternoon, where the max temperature
was at 5m.  The temperature differences were about 0.2 C from 0.5m to 5m and
back from 5m to 20m.  I believe the temperature sensors by now.  Later this
evening, it broke down and became completely stable throughout the profile.

72: Temp_RH, Site all stns, Sat 08-Dec-2001 13:59:34 MST, TRH 703
We now have in the trailer two "spare" T/RHs.  One is 704 which I just removed
from the tower at 20m due to temperature readings that were 0.2 C low (the 
profile looks MUCH better now).

The other is 703 which Steve had removed from 2m a while back.  Steve said that
the fan was bad.  I took it apart and found that the ribbon cable was cut.  I
spliced in a section of normal cable, but noticed some bare insulation on the
wires directly connected to the fan.  I didn't patch these.  Thus, 703 should
now be considered a working spare, though it should have a better repair job
sometime.


70: Status, Site all stns, Sat 08-Dec-2001 10:12:55 MST, serial numbers and heights
These are the serial numbers for sensors that I know of:

Main:	TRH	Prop	CSAT	KH20	Baro	Logger
20	704	0007	0540	1394
15	701	0006
10	201	0003	0537	1393
5	001	0002	
2	004	0001	0536	
1	702	0004
0.5	005
Bare:
2	007		0539	1392	1002	1

Note that I'm just about to change the TRH at 20m to SN 104. (Done 10:55.)

Here are the heights of these sensors:

TRH
20	20.40+0.31-0.04-0.57 = 20.10
15	16.72-0.72-0.04-0.57 = 15.39
10	11.20-0.67-0.04-0.57 = 9.92
5	5.78+0.31-0.04-0.57 = 5.48
2	2.10+0.46-0.04-0.57 = 1.95
1	2.10-0.55-0.04-0.57 = 0.94
0.5	2.10-1.02-0.04-0.57 = 0.47

Prop
20	20.40-0.84-0.06+0.45 = 19.95
15	14.88-0.01+0.45 = 15.32
10	9.46+0.05-0.06+0.45 = 9.90
5	5.78-0.86-0.06+0.45 = 5.31
2	2.10-0.51-0.06+0.45 = 1.98
1	2.10-1.57-0.06+0.45 = 0.92

CSAT
20	20.40-0.89-0.06+0.05 = 19.50
10	11.20-0.89-0.06+0.05 = 10.30
2	2.10+0.04+0.05 = 2.19

KH20 relative to CSAT: 20cm down, 35cm back
LI6251 inlet frit relative to CSAT: 10cm down, 33cm back
LI7500 center of path relative to CSAT: 40cm down, 67cm back





		20.40-0.84-0.06+0.45
59: Radiometers, Site all stns, Thu 06-Dec-2001 12:25:28 MST, radiometers and soil sensors
Radiometers and soil sensors

Radiation stands were set up at the Main site and the Bare site on 28 Dec''01

The radiometers were installed at Bare on 29 Nov''01
The radiometers and snow sensor were installed at Main on 30 Nov''01

The soil sensors were installed at Main on 30 Nov''01
The soil sensors were installed at Bare on 30 Nov''01

The radiometers were leveled and fan function were checked at Main on 01 Dec''01
The radiometers were leveled and fan function were checked at Bareon 01 Dec''01

The complement of sensors is shown in ...../doc/radioms.idr (to be copied from
~/deploymnets/floss02/radioms.idr)

44: Barometer, Site all stns, Tue 04-Dec-2001 15:49:34 MST, Pressure error
I've been puzzled by the pressure difference between main and bare.  The height
difference only justifies about 0.3mb (bare lower).  (The difference in height
from the topos was about 13' = 4m at ~11m/mb.)  So, I just made a difference
plot and noticed that it varied like temperature by 2mb over 15C.  A 
scatterplot of the pressure difference vs. temperature (see attached) has a 
perfect correlation.  This sounds like the same problem we saw in OASIS98.  
We should get rid of the non-compensated sensor.
43: Soil, Site all stns, Tue 04-Dec-2001 14:39:08 MST, soil moisture
I made some hand-held TRIME measurements near the soil probes at the main tower
today.  At two sites I got: 7.3% and 6.9%.  Note that the first readings were
both 6.9%.  The first site showed increasing values over the first minute.

These were done about 10:15 this morning.
 
Later in the day, (4:15), took readings at bare site: 4.2%, 4.7%, and 5.0%.
I think that the 4.7 was about like the 615 probe.


32: , Site all stns, Mon 03-Dec-2001 10:51:29 MST, task list
Here are the tasks I can think of:

Daily:
- check_aster
- daily status to logbook
- snow photograph
- trime soil moisture reading
- read cal gas gauges

End of Walden visits:
- swap backup disk
29: Status, Site all stns, Mon 03-Dec-2001 08:46:28 MST, daily status
Daily status:

Weather: almost completely clear (some clouds over mountains), snow disappearing
(10-20% bare?).  I'll try to take tower photos today.

cosmos-up
marigold-up (goldenrod hasn't crashed since replaced)
stn2-up
aster-gives RPC error

main site:
props    all okay
T/RH     something has been weird at 15 and/or 20m.  Maybe I'll just swap these.
Tc       5.75m died.  Others okay?  I'll have to make a daily plot that adds in
         the reference.  The 21x values appear to be somewhat noisy.
sonics   okay, though more spikes than I'd like (20-30 overnight at all levels)
kryptons okay, the 2m sensors have had spikes (large variance) in the data.
         I'll clean these soon, though the spikes are probably from blowing snow
licorCO2 okay, cal gases are cycling
rad	 okay I guess.  Haven't looked at daily plots yet.
soil	 okay I guess.  "
icedet   no action (which is good)
levels   okay
baro     okay
precip   okay?  reading zero.  I should do a calibration.
snowdepth okay?  may be seeing blowing grasses
power	 no problems

bare site:
T/RH	 okay?  The fan acted strange last night, perhaps due to power
sonic	 good.  The fewest spikes (perhaps the least blowing snow?)
krypton  okay.  weird values on cockpit probably from my work last night
heater   false high readings.  need to investigate this.
rad      okay.  need to look at daily plots
soil     okay.  "
levels   weird excursions.  need to check this
baro     okay
gps      okay
power    mostly only charging during day -- problem with wind generator?

need to rabbit-proof all cables!




26: Communications, Site all stns, Sun 02-Dec-2001 14:47:55 MST, data network configuration
Data Network

BCAP: Breezecom Access point in base trailer
BCSA: Breezecom Station Adaptors at tower and in OSU box

aster: NCAR base system, Dell Latitude, linux
tiny:  OSU linux system
ambra: Windows 486 running Breezecom DS.11 diagnostic software

****************** 192.168.10.0 *************************************

aster -- 5port --  BCAP ---RF---  BCSA --cable-- 5port ----- cosmos
ambra -- hub        |             tower          hub   ----- marigold
tiny  --            |                                  ----- web cam
                    RF
                    |
                    |-------RF--  BCSA --cable-- 4port ----  cr5000
		                  OSU            hub   ----  lunch

*********************************************************************

These addresses are assigned on the Data Network:

  192.168.10.1	aster
  192.168.10.10-20  allocated dynamically by DHCP server on aster
		    Breezecom radios and Ambra PC

NCAR Data Acq Systems
  192.168.10.30	cosmos	
  192.168.10.31	cosmos-bp	
  192.168.10.32	goldenrod	
  192.168.10.33	marigold	
  192.168.10.34	marigold-bp	
  192.168.10.35	arnica	

OSU PCs
  192.168.10.50 lunch
  192.168.10.51 cr5000




25: Status, Site all stns, Sun 02-Dec-2001 08:33:21 MST, Dec 2 morning status
Weather:
  light snow last night, windy.  Small 8" drifts on refuge road.


Data system:
  up.  Marigold/arnica made it through the night!

Sonics:
  
  some non-zero diag values on all sonics. A few values >=16 which
    indicate a lost sample.
  all sonics also show a few scattered non-zero flags

  heater value at bare site always on

Kryptons: OK
  
TRH: OK

Props: OK

Thermocouples:  Sean had 4 deployed last night.  2 of the 4 are no
  longer reporting, assumed broken.

Radiation: values appear to be in acceptable range - did not do
  a thorough check.  
  Xlevel.bare and Ylevel.bare show some values around 10.

PAM power system:  still some question about whether wind generator
  voltage is getting to charging system.
  Current values: 
	Vbatt: 12.4 V
 	Icharge: 1.5 - 2.6 A  (sun through thin clouds, 7 m/s winds)
	Iload: 1.7 A
	

22: Status, Site all stns, Sat 01-Dec-2001 16:23:38 MST, Operational!

All NCAR sensors deployed and reporting data, however quality
checks have been minimal.

Prop angles have been entered, so prop wind directions should be
correct.

Licor co2 is deployed.  See other entry about calibration gas sequence.

Sean Burns has installed 4 (out of 20) thermocouples on the tower.
They are at 14, 15.5, 17, and 18.5 meters.

Kurt, SteveS and Tony left at about 3:00 this afternoon.  SteveO due
back tomorrow.


15: soil, Site all stns, Sat 05-Aug-2000 08:53:54 PDT, Soil water gravimetric data
FLOSS01 Soil Gravimetric water fraction. 
All the soil moisture data is to be entered in this logbook file


Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.  
 62 SO  06 Dec 01 Bare    0-2   A1 18.8    16.5    2.3    0.14
                                A2 19.6    17.4    2.2    0.13  = 0.14
                          2-5   B1 23.8    20.8    3.0    0.14
                                B2 24.7    21.7    3.0    0.14  = 0.14
                          5-8   C1 34.8    31.7    3.1    0.10
                                C3 36.4    33.7    2.7    0.08  = 0.09
                          8-10+ D1 30.9    28.0    2.9    0.10
                                A3 30.6    28.2    2.4    0.09 	= 0.10 	                                      
---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
 68 SO  08 Dec 01 tower   0-3   B1 11.8     8.1    3.7    0.05
                                B2 12.3     6.6    5.7    0.86  = ?
                          3-6   C1 32.9    29.4    3.5    0.12
                                C3 32.0    28.1    3.9    0.14  = 0.13
                          6-8+  D1 40.8    37.7    3.1    0.08
                                A3 36.1    33.3    2.8    0.08  = 0.08

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
 71 SO  08 Dec 01 tower 0-1.5   A1 13.3     2.9   10.4    3.6
                                A2 10.9     2.4    8.5    3.5   = 3.6
                        1.5-4.5 B1 27.0    12.2   14.8    1.2
                                B2 15.4     9.7    5.7    0.59  = 0.90
                        4.5-7.5 C1 47.9    43.1    4.8    0.11
                                C3 33.8    30.5    3.3    0.11  = 0.11
                        7.5-9.5 D1 33.8    31.7    2.1    0.07
                                A3 34.7    32.8    1.9    0.06  = 0.07

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
118 SS  17 Jan 02 tower   0-5   A1 14.3    15.9
                                A2 16.3    17.4                     ? 
                                A3 16.3    18.0


---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
118 SS  17 Jan 02 sage    0-5   B1 29.7    25.7   4.0     0.16
                                B2 31.0    26.4   4.6     0.17   = 0.16
                                B3 31.3    27.0   4.3     0.16

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
141 SO  18 Feb 02 tower   0-9.5 A1 36.3    33.3    3.0    0.09
                                A2 26.7    24.3    2.4    0.10   = 0.11
                                A3 28.8    25.4    3.4    0.13

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
146 TD  20 Feb 02 sage    0-5   A1 19.4    16.5    2.9    0.18
                                A2 20.4    17.4    3.0    0.17  = 0.18
                                A3 23.4    19.7    3.7    0.19

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
146 TD  20 Feb 02 bare    0-5   B1 19.6    16.7    2.9    0.17     
                                B2 23.9    20.1    3.8    0.19  = 0.18
                                B3 27.1    22.7    4.4    0.19

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
146 TD  20 Feb 02 tower   0-5   C1 17.6    13.2    4.4    0.33
                                C2 17.0    12.7    4.3    0.34  = 0.35
                                C3 16.6    12.0    4.6    0.38

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
147 TD  22 Feb 02 tower   0-2   D1  5.8     3.6    2.2    0.61
                                D2  4.7     2.8    1.9    0.68  = 0.64
                                D3  5.9     3.6    2.3    0.64
                          2-4   E1 14.9    11.9    3.0    0.25
                                E2 17.3    13.7    3.6    0.26  = 0.25
                                E3 19.0    15.2    3.8    0.25
                          4-8   F1 19.2    17.1    2.1    0.12
                                F2 18.7    16.7    2.0    0.12  = 0.12
                                F3 20.4    18.3    2.1    0.11
                          8-16  G1 26.3    24.9    1.4    0.06
                                G2 23.0    21.8    1.2    0.06  = 0.06
                                G3 29.0    27.4    1.6    0.06

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
157 TD  27 Feb 02 Sage    0-2   A1 22.1    18.1
                                A2 25.9    21.2
                                A3 26.1    21.5

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
157 TD  27 Feb 02 Sage    2-5   B1 23.8    21.5
                                B2 21.9    19.8
                                B3 26.1    20.6

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
157 TD 27 Feb 02  Sage    5-10  C1 30.7    27.9
                                C2 29.6    26.8
                                C3 22.9    20.6

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
157 TD 27 Feb 02  Bare    0-2   D1 11.3     9.5
                                D2  8.9     7.6
                                D3 11.0     9.3

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
157 TD 27 Feb 02  Bare    2-5   E1 21.7    18.6
                                E2 19.9    16.8
                                E3 19.8    16.9

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
157 TD 27 Feb 02  Bare    5-10  F1 24.2    21.4
                                F2 24.7    21.8
                                F3 24.5    21.7

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
162 TD  28 Feb    Tower   0-2   A1 3.9      2.0
                                A2 5.1      2.5
                                A3 3.9      1.9

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
162  TD  28 Feb    Tower   2-5   B1 7.0      5.7
                                 B2 9.1      7.6
                                 B3 7.2      6.1

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
162 TD  28 Feb    Tower   5-10  C1 24.5    23.2
                                C2 27.9    26.3
                                C3 27.6    26.0

---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
171     13 Mar    Tower   0-2   IA   9.2    3.5
                                IB   9.6    3.6

                          2-5  IIA  24.6   16.9
                               IIB  31.6   23.2
                                
                         5-10 IIIA  40.2   36.7
                              IIIB  39.2   35.9

Oven-dried aliquots weighed on 26 Mar 02. See note in 171
---------------------------------------------------------------------------
Lgbk #  date	  Site    Depth    Fresh   Dry    Diff  Grav.Frac.
177 TD  26 Mar    Tower   0-2  IA   8.7   4.2
                               IB  10.0   4.6
                               IC  11.0   4.9

                          2-5 IIA  29.3  22.9
                              IIB  35.1  27.2
                              IIC  31.8  23.9

                         5-10 IIIA 30.2  25.2
                              IIIB 36.6  29.5
                              IIIC 34.1  28.1
Oven-dried aliquots weighed on April 2. GDM





















14: Status, Site all stns, Thu 29-Nov-2001 17:49:08 MST, status
Today's work:
- installed all T/RHs on main
- intalled all props on main
- installed 2m CSAT & KH20 on main
- set up rad stand (but not installed in final position)
- set up bare site infrastructure and some sensors

Other:
- Larry Mahrt and Dean Vickers arrived, staked out their sites and organized their equipment
- Sean Burns arrived, will set up tomorrow
- Mark Lanier stopped by for a visit

8: Adams, Site all stns, Tue 20-Nov-2001 16:15:06 GMT, manual control note
The ttl manual control is reverse logic.
So ttl_on turns power off and ttl_off turns power on.
I confirmed this on marigold. I assume it is the same
for cosmos.
5: Communications, Site all stns, Sat 10-Nov-2001 23:44:39 GMT, Breezecom settings
AP: Access point, this will be the radio at the base, in the trailer
SA: Station adaptor, this is the radio out at the tower.

Community strings:
  read: ncar/atds
  write: (see gordon)

  WLAN parameters -> ESS ID: see gordon
  WLAN parameters -> fixed channel, #11

For testing in the building, I have 2 little dipole antennals
on each radio:
  WLAN parameters -> Transmit Diversity -> both antennas

Security:
  Default key ID: First
  WEP Keys: see gordon
    1st: 
    2nd: 
    3rd: 
    4th: 


Breezecoms were showing a high FCS error rate.  According to the
DS11 manual this is a CRC error statistic.  The adams seemed to
be suffering, losing data, and crashing.

Tried various things, which didn't work, including 1/2 meter dipole antenna
on AP, moving radios further apart, resetting to factory defaults.

What did help was changing the channel number from 6 to 11.  It appears
that another wireless ethernet in the building is using channel 6.

According to the breezecom manual (Appendix, page B-2), only channels
1,6 and 11 are non-overlapping.

Settings:
  SA: WLAN Parameters -> Channel Settings -> Fixed Channel, #11
  AP: WLAN Parameters -> Channel 11

To view the FCS statistic, select the SA, click on the "Counters" tab.
A useful tool is to click on the FCS number and drag it to the
graph image.  Once the graph appears, click on scale to change
the scale to 0-10 counts. Currently we're seeing an FCS error 
rate of 5%, with one or two errors every 5 seconds.  This 
seems to be acceptable.


4: Communications, Site all stns, Sat 10-Nov-2001 22:44:47 GMT, Adam/network testing

The adams have been transmitting via the breezecoms since yesterday
(fri Nov 9) afternoon.  The security was set to "secured key",
which SCD recommends.

The network has not been working well since then.
The adams would start reporting lost samples
and crashed several times.

The radios are set up with the indoor dipole antennas.  Until now
I had both antennas connected.   The DS11 program was showing a bunch
of errors.  The FCS Error rate was about 75%.

According to the DS11 manual, only antenna #1 should be connected.
After making that change, things have improved. We'll see how it goes.
The breezecoms are about 8 feet apart in the staging area.

Current error status shown on DS11 program, 

  SA showing 80-85 % signal quality,  RSSI -27dBm (62)
     FCS Error 11%

  AP: FCS Error 11%

It looks like the FCS error rate is a critical number.